OpenStudy (barrelracer011):

Will medal :) Would this be a positive or negative slope on a graph???? 15x - 5y = 30.

OpenStudy (anonymous):

Negative I think, positive+positive = negative.

ganeshie8 (ganeshie8):

\(Ax+By = C\) slope = \(\large -\frac{A}{B}\)

OpenStudy (barrelracer011):

OK thank you it helps alot

ganeshie8 (ganeshie8):

np :) or another way is to convert the given equation into 'slope-intercept' form

ganeshie8 (ganeshie8):

15x - 5y = 30 subtract 15x both sides -5y = -15x + 30 divide -5 both sides y = 3x - 6

OpenStudy (barrelracer011):

so it would be -6 over 2 or -2 or 6

ganeshie8 (ganeshie8):

compare it wid the slope-intercept form :- y = mx + b

ganeshie8 (ganeshie8):

m = 3 so, slope = 3

OpenStudy (anonymous):

thx for the medal